# An Evaluation of Sakana’s AI Scientist for Autonomous Research: Wishful Thinking or an Emerging Reality Towards ‘Artificial General Research Intelligence’ (AGRI)?

### v1

@@ 2.7. Costs and Effort
When assessing the AI Scientist’s potential, cost is a crucial factor — and surprisingly, it remains low. The total cost for our main experiments, which included generating 10 ideas, running experiments for 7 ideas (from 12 total, including 2 seed ideas, of which 5 failed), and producing 7 manuscripts with reviews, amounted to only \$42 USD — an average of \$6 per manuscript. Given that our experiments were less complex than typical recommender systems research or those conducted by Sakana, their claim that a full research paper can be produced for around \$15 seems realistic.
However, human labor is still required. Setting up the AI Scientist took approximately 5 hours, while implementing the experimental template required an additional 15 hours. It must be noted that if we were to re-implement similar pipelines, the required time would decrease due to the learning curve. If we were to implement more complex experiments, time would increase. Beyond this, we spent about 10 more hours brainstorming initial seed ideas, reviewing outputs, and refining results. Excluding the initial setup time, this totals 25 hours of human effort —approximately 3.5 hours per manuscript — mainly from an undergraduate student.
@@ 2.8. Limitations
Our own evaluation had limitations: we used a single dataset (MovieLens-100k) and one research domain (Green Recommender Systems), relied on fixed seed ideas and user-defined templates, and required manual adaptations for hardware compatibility and troubleshooting. While these constraints ensured a controlled assessment, they may affect generalizability and reproducibility. Despite this, we believe our findings remain meaningful.
@@ 3.1. Researchers should familiarize themselves with research-assistive AI tools
At present, research-assistive tools like the AI Scientist hold little practical value. In contrast, more general AI tools, such as ChatGPT for text generation and GitHub Copilot for programming, already provide substantial benefits in research, improving writing, coding, and brainstorming efficiency. Researchers who ignore these readily available AI-powered assistants are already at a disadvantage compared to their peers who integrate them into their workflows.
@@
Early-career researchers, in particular, must closely monitor these developments. The use of AI in research will soon become indispensable. Those who fail to adopt AI tools risk falling behind their peers, akin to a researcher today relying on a typewriter and calculator instead of adopting a computer with spreadsheet and programming language support. The future of research belongs to those who embrace AI-enhanced workflows and integrate them into their daily scientific practice.
@@ 4. Discussion and Conclusion
In summary, while the AI Scientist falls short of its grand claims, it offers a glimpse into the future of AI-driven scientific discovery and towards ‘Artificial General Research Intelligence’ (AGRI). With responsible development and integration, AI could enhance knowledge generation, improve reproducibility, and streamline the research process in ways we are only beginning to explore. However, realizing this potential requires careful consideration of AI’s limitations, ethical challenges, and long-term implications for the scientific community. The time to act and to participate in evaluating, exploring, discussing and developing AI-driven research agents is now!

### v2

@@ 2.7. Costs and Effort
When assessing the AI Scientist’s potential, cost is a crucial factor — and surprisingly, it remains low. The total cost [^20] We report only costs of OpenAI’s API; other costs such as electricity for our laptop or university cluster are ignored for our main experiments, which included generating 10 ideas, running experiments for 7 ideas (from 12 total, including 2 seed ideas, of which 5 failed), and producing 7 manuscripts with reviews, amounted to only \$42 USD — an average of \$6 per manuscript. Given that our experiments were less complex than typical recommender systems research or those conducted by Sakana, their claim that a full research paper can be produced for around \$15 seems realistic.
However, human labor is still required. Setting up the AI Scientist took approximately 5 hours, while implementing the experimental template required an additional 15 hours. It must be noted that if we were to re-implement similar pipelines, the required time would decrease due to the learning curve. If we were to implement more complex experiments, however, time would increase. Beyond this, we spent about 10 more hours brainstorming initial seed ideas, reviewing outputs, and refining results. Excluding the initial setup time, this totals 25 hours of human effort — approximately 3.5 hours per manuscript — mainly from an undergraduate student.
@@ 2.8. Limitations
Our own evaluation had limitations: we used a single dataset (MovieLens-100k) and one research domain (Green Recommender Systems), relied on fixed seed ideas and one template. While these constraints ensured a controlled assessment, they may affect generalizability and reproducibility. Despite this, we believe our findings remain meaningful.
@@ 3.1. Researchers should familiarize themselves with research-assistive AI tools
At present, tools like the AI Scientist hold little practical value. In contrast, more general AI tools, such as ChatGPT for text generation and GitHub Copilot for programming, already provide substantial benefits in research, improving writing, coding, and brainstorming efficiency. Researchers who ignore these readily available AI-powered assistants are already at a disadvantage compared to their peers who integrate them into their workflows.
@@
Early-career researchers, in particular, must closely monitor these developments. The use of AI in research will soon become indispensable. Those who fail to adopt AI tools risk falling behind their peers, akin to a researcher today relying on a typewriter and calculator instead of adopting a computer with text editor and programming language support. The future of research belongs to those who embrace AI-enhanced workflows and integrate them into their daily scientific practice.
@@ 4. Discussion and Conclusion
In summary, while the AI Scientist falls short of its grand claims, it offers a glimpse into the future of AI-driven scientific discovery and towards ‘Artificial Research Intelligence’ (ARI). With responsible development and integration, AI could enhance knowledge generation, improve reproducibility, and streamline the research process in ways we are only beginning to explore. However, realizing this potential requires careful consideration of AI’s limitations, ethical challenges, and long-term implications for the scientific community. The time to act and to participate in evaluating, exploring, discussing and developing AI-driven research agents is now!